# Dual-VQE: A quantum algorithm to lower bound the ground-state energy

### v2

@@ Abstract
The variational quantum eigensolver (VQE) is a hybrid quantum–classical variational algorithm that produces an upper-bound estimate of the ground-state energy of a Hamiltonian. As quantum computers become more powerful and go beyond the reach of classical brute-force simulation, it is important to assess the quality of solutions produced by them. Here we propose a dual variational quantum eigensolver (dual-VQE) that produces a lower-bound estimate of the ground-state energy. As such, VQE and dual-VQE can serve as quality checks on their solutions; in the ideal case, the VQE upper bound and the dual-VQE lower bound form an interval containing the true optimal value of the ground-state energy. The idea behind dual-VQE is to employ semi-definite programming duality to rewrite the ground-state optimization problem as a constrained maximization problem, which itself can be bounded from below by an unconstrained optimization problem to be solved by a variational quantum algorithm. When using a convex combination ansatz in conjunction with a classical generative model, the quantum computational resources needed to evaluate the objective function of dual-VQE are no greater than those needed for that of VQE. We also show that the problem is well suited for classical pretraining using matrix product states and these methods help warm-start the optimization. We simulated the performance of dual-VQE on the transverse-field Ising model with and without pretraining and found that, for the example considered, while dual-VQE training is slower and noisier than VQE, it approaches the true value with an error of order $10^{-2}$.
@@ I.1 Background
where $\mathcal{D}$ is the set of density matrices (positive semi-definite matrices with trace equal to one). Since the objective function $\operatorname{Tr}[H\rho]$ is linear in $\rho$ and the set of density matrices is convex, the optimal value of (1) is obtained on the boundary of the set $\mathcal{D}$ (i.e., the set of pure states). It then follows that
@@
where $\mathcal{P}$ is the set of state vectors. In a variational approach, the set of pure states is approximated by a parameterized set $\{|\psi(\theta)\rangle\}_{\theta\in\Theta}$ of states, where $\Theta$ is a parameter set and $\theta$ is a parameter vector. The equality in (2) is preserved if the set of parameterized pure states contains an optimal pure state. In reality, guaranteeing this condition is not possible, leading to the following upper bound on the ground-state energy:
@@
An example of wide interest is when $n\in\mathbb{N}$ and $H$ is a $2^{n}\times 2^{n}$ Hermitian matrix, describing the local interactions that take place between $n$ spin-1/2 particles. In this case, $H=\sum_{j=1}^{L}H_{j}$, where each $H_{j}$ is a Hermitian matrix of constant size that acts non-trivially on a constant number of particles. Observing that the optimization in (1) is a semi-definite program (SDP) [5], one could attempt to solve it by means of well known approaches used to solve SDPs; however, the computational complexity of such approaches is exponential in $n$ and thus infeasible. Standard classical approaches for solving (1) in the general case suffer from the same problem, or they avoid it by not allowing for the optimization to include highly entangled states (see [6] for a review of some of these methods).
VQE instead parameterizes the set of state vectors by a parameterized quantum circuit $U(\theta)$ acting on an easily preparable initial state $|\psi_{0}\rangle$, leading to $|\psi(\theta)\rangle\coloneqq U(\theta)|\psi_{0}\rangle$, and measures the expectation of each observable $H_{j}$ with respect to $|\psi(\theta)\rangle$ in order to obtain an estimate $\widehat{E_{j}}$ of $\langle\psi(\theta)|H_{j}|\psi(\theta)\rangle$. Then the sum $\sum_{j}\widehat{E_{j}}$ is an estimate of the desired expectation $\langle\psi(\theta)|H|\psi(\theta)\rangle$. The estimates of $\langle\psi(\theta)|H|\psi(\theta)\rangle$ and its gradient can be fed to a classical optimization procedure, and this process iterated until convergence or a maximum number of iterations is reached. As a consequence of (3), this procedure in the ideal case produces an estimate of an upper bound on the ground-state energy $\lambda_{\min}(H)$.
@@ I.2 Main contribution: Dual-VQE method
The dual-VQE method involves 1) converting inequality constraints to equality constraints by means of positive semi-definite slack variables, 2) converting a constrained optimization into an unconstrained one by means of a penalty term in the objective function, 3) replacing positive semi-definite slack variables with scaled parameterized quantum states that are efficiently preparable on quantum computers, and 4) estimating various terms in the objective function by means of the destructive swap test [7] and sampling estimates of expectations of observables. Our companion paper [8] explores this general approach in a much broader context beyond the ground-state energy problem.
Although there have been various classical algorithms proposed for obtaining lower bounds on the ground-state energy [9, 10, 11, 12, 13], they all ultimately suffer from the exponential increase of the space on which $H$ acts, as the number of particles increases. Our approach is fundamentally different, involving a reformulation of the original optimization problem in terms of its SDP dual, and then bounding that quantity from below using a variational quantum algorithm. Our approach is also complementary to [14, Section VI-A], which does not take a variational approach, nor does it provably obtain a lower bound on the ground-state energy.
@@ I.3 Pretraining with matrix product states
A key issue that plagues variational quantum algorithms is the presence of barren plateaus in the training landscape [15, 16, 17], meaning that the gradient (or equivalently cost differences [18]) vanishes exponentially fast in the number of qubits. At their core, barren plateaus arise from a curse of dimensionality [19] stemming from the exponentially scaling size of the Hilbert space. On the other hand, standard techniques for provably avoiding barren plateaus do so by encoding the relevant dynamics in a smaller subspace making it effectively possible to classically simulate the parameterised quantum circuit [19].
@@ II Theoretical derivation
where $\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}$ is the SDP dual of (1). To see this, recall that the matrix inequality $A\geq B$ holds for Hermitian matrices $A$ and $B$ if ABA-B is a positive semi-definite matrix. Then consider that
@@
Now that we have recalled this standard result, let us focus on showing how to estimate the left-hand side of (4) on a quantum computer. Recalling that the matrix inequality $\eta I\leq H$ is equivalent to the existence of a positive semi-definite matrix $W$ such that $H-\eta I=W$, we conclude that
@@
Our next observation, related to observations used in [23], is that every positive semi-definite matrix $W\neq 0$ can be written as a scaled density matrix, i.e., $W=\nu\omega$, where $\nu=\operatorname{Tr}[W]$ and $\omega=W/\nu\in\mathcal{D}$, implying that
@@
where $c>0$ is a penalty parameter and $\left\|Z\right\|_{2}\coloneqq\sqrt{\operatorname{Tr}[Z^{{\dagger}}Z]}$ is the Hilbert–Schmidt norm of a matrix $Z$. Note that the equality in (13) holds whenever the penalty term is faithful (specifically, we invoke [24, Proposition 5.2.1] to justify this conclusion, setting $\lambda_{k}=0$ therein for all $k$). By faithfulness, we mean that the penalty term is equal to zero if and only if $H-\eta I=\nu\omega$. This is true for the Hilbert–Schmidt norm, and this particular norm has the convenient advantage that it can be estimated by quantum sampling procedures that we recall below.
The final step is to modify the optimization in (13) to be over a parameterized set $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of density matrices instead of being over all density matrices, which implies that
@@
The function $f^{\ast}(c)$ in (15) provides a lower bound on the ground-state energy of the Hamiltonian $H$ in the limit $c\to\infty$; i.e.,
@@
In the rest of the work, we use the term “objective function” to refer to $\eta$, and “penalty” to refer to the sum of all terms in (16) that are multiplied by the penalty parameter $c$, i.e.,
$\operatorname{Tr}[H^{2}]+\eta^{2}2^{n}+\nu^{2}\operatorname{Tr}[\omega(\theta)^{2}]\\ -2\eta\operatorname{Tr}[H]-2\nu\operatorname{Tr}[H\omega(\theta)]+2\eta\nu.$ (18)
@@ III Algorithm construction
The basic idea behind the dual-VQE algorithm is to follow the standard penalty method approximately [24, Section 5.2]: define a sequence $\left(f^{\ast}(c_{k})\right)_{k\in\mathbb{N}}$ of optimization problems where $\left(c_{k}\right)_{k\in\mathbb{N}}$ is a monotone increasing sequence of penalty parameters. We then solve the optimization for each $k$ and use the solution as the initial guess for the next iteration. Under the assumption that each optimization $f^{\ast}(c_{k})$ can be solved, the method is guaranteed to converge to the correct solution. In our simulations, we use a variety of update methods to change the value of $c$ as the optimization progresses. In some cases, we pick $c$ to be a high value and keep it constant throughout the optimization. In other cases, we begin with a smaller value, and ramp it up as the iteration count grows.
@@
As is common in Hamiltonian models relevant for physics and chemistry, the number of non-zero coefficients in the tuple $\left(\alpha_{\overrightarrow{x}}\right)_{\overrightarrow{x}}$ is polynomial inn~n, so that the objective function in (16) can be efficiently estimated. In this work, we focus on the transverse-field Ising model given in (32), which falls under this category.
@@
III.1 Parameterization
There are at least two general ways of parameterizing the set of mixed states, called the purification ansatz [25, 23, 26] and the convex combination ansatz [27, 28, 26, 29] (see also [8]). Focusing on the first one to start, it is based on the purification principle [30, 31, 32], in which every mixed state is realized as the marginal of a pure state on a larger system. For the purification ansatz, we simply start with a parameterized family $\left\{|\psi(\theta)\rangle_{RS}\right\}_{\theta\in\Theta}$ of pure states of a reference system $R$ and the system $S$ of interest, and then we obtain a parameterized family $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of mixed states via the partial trace $\omega(\theta)=\operatorname{Tr}_{R}[|\psi(\theta)\rangle\!\langle\psi(\theta)|_{RS}]$. We can estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ by means of the destructive swap test [7] (see, e.g., [33, Section 2.2] for a precise statement of the algorithm). We can estimate the gradient of the objective function in (17) by means of the parameter-shift rule [34, 35, 36], simply because
@@ III.1 Parameterization
$F$ is the unitary and Hermitian swap operator, and we have suppressed some system labels in (24) for brevity (see also [26, Eq. (C2)] in this context).
The convex combination ansatz is based on the fact that every mixed state can be written as a convex combination of pure orthogonal states. Following this idea, we obtain a parameterized family $\left\{\omega(\varphi,\gamma)\right\}_{\varphi,\gamma\in\Theta}$ by separately parameterizing the eigenvalues and eigenvectors as follows:
@@
where $\left(p_{\varphi}(x)\right)_{x}$ is a parameterized probability distribution, $U(\gamma)$ is a parameterized unitary, and $\left\{|x\rangle\right\}_{x}$ is the standard computational basis. The parameterized probability distribution $p_{\varphi}$ can be realized by a neural-network-based generative model [37, 38, 39] or a quantum circuit Born machine [40], while the parameterized unitary $U(\gamma)$ can be realized by a parameterized circuit.
Interestingly, although it is necessary to use a quantum computer to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ for the purification ansatz, it is not necessary to do so when using the convex combination ansatz in conjunction with a neural-network-based generative model: it suffices to employ a classical sampling approach, thus further reducing the quantum resource requirements for dual-VQE (see also [26, Eqs. (20)–(23)] in this context). This is because $\operatorname{Tr}[\left(\omega(\varphi,\gamma)\right)^{2}]=\sum_{x}p_{\varphi}^{2}(x)$, so that one can repeatedly sample from $p_{\varphi}$ and perform a collision test to estimate $\sum_{x}p_{\varphi}^{2}(x)$. Thus, for the convex combination ansatz used in the aforementioned way, one only needs a quantum computer to estimate the expectation
@@ III.2 Pretraining
Pretraining broadly refers to a set of relatively computational inexpensive techniques that kick start (or warm start) an optimization process, as opposed to a random start. In several domains, pretraining has been shown to be advantageous in reducing the training time and accelerating convergence [42, 43].
In this section, we describe the details of the particular pretraining process used in this work. The overall process consists of two phases: pretraining, followed by a translation into a quantum circuit. The pretraining is conducted using matrix product states (MPS) [44], and the final translation consists of a combination of two techniques – analytic decomposition (AD) and optimizing decomposition (OD), as proposed in [45, 46]. Important preliminaries on tensor networks and matrix product states can be found in Appendix A, and key ideas of the pretraining process, including an in-depth review of AD and OD, can be found in Appendix B.
Before delving into the specifics of the algorithm, we aim to establish the rationale behind pretraining using MPS. Our motivation for this approach stems from Refs. [47, 21], which explored solving the Variational Quantum Eigensolver (VQE) problem by pretraining it with MPS. The study demonstrated that parameterized quantum circuits for VQE exhibited significant performance improvements when used in conjunction with a pretrained initialization, in contrast to a random initialization. When applied to the dual-VQE problem, we find that pretraining provides a significant kick start, with the advantage increasing with maximum bond dimension of the MPS.

### v3

# Dual variational quantum eigensolver: A quantum algorithm to lower bound the ground-state energy

@@ Abstract
The variational quantum eigensolver (VQE) is a hybrid quantum–classical variational algorithm that produces an upper-bound estimate of the ground-state energy of a Hamiltonian. As quantum computers become more powerful and go beyond the reach of classical brute-force simulation, it is important to assess the quality of solutions produced by them. Here we propose a dual variational quantum eigensolver (dual-VQE) that produces a lower-bound estimate of the ground-state energy. As such, VQE and dual-VQE can serve as quality checks on their solutions; in the ideal case, the VQE upper bound and the dual-VQE lower bound form an interval containing the true optimal value of the ground-state energy. The idea behind dual-VQE is to employ semidefinite programming duality to rewrite the ground-state optimization problem as a constrained maximization problem, which itself can be bounded from below by an unconstrained optimization problem to be solved by a variational quantum algorithm. When using a convex combination ansatz in conjunction with a classical generative model, the quantum computational resources needed to evaluate the objective function of dual-VQE are no greater than those needed for that of VQE. We also show that the problem is well suited for classical pretraining using matrix product states and these methods help warm-start the optimization. We simulated the performance of dual-VQE on the transverse-field Ising model with and without pretraining and found that, for the example considered, while dual-VQE training is slower and noisier than VQE, it approaches the true value with an error of order $10^{-2}$.
@@ I.1 Background
where $\mathcal{D}$ is the set of density matrices (positive semidefinite matrices with trace equal to one). Since the objective function $\operatorname{Tr}[H\rho]$ is linear in $\rho$ and the set of density matrices is convex, the optimal value of (1) is obtained on the boundary of the set $\mathcal{D}$ (i.e., the set of pure states). It then follows that
@@
where $\mathcal{P}$ is the set of state vectors. In a variational approach, the set of pure states is approximated by a parametrized set $\{|\psi(\theta)\rangle\}_{\theta\in\Theta}$ of states, where $\Theta$ is a parameter set and $\theta$ is a parameter vector. The equality in (2) is preserved if the set of parametrized pure states contains an optimal pure state. In reality, guaranteeing this condition is not possible, leading to the following upper bound on the ground-state energy:
@@
An example of wide interest is when $n\in\mathbb{N}$ and $H$ is a $2^{n}\times 2^{n}$ Hermitian matrix, describing the local interactions that take place between $n$ spin-1/2 particles. In this case, $H=\sum_{j=1}^{L}H_{j}$, where each $H_{j}$ is a Hermitian matrix of constant size that acts nontrivially on a constant number of particles. Observing that the optimization in (1) is a semidefinite program (SDP) [49], one could attempt to solve it by means of well known approaches used to solve SDPs; however, the computational complexity of such approaches is exponential in $n$ and thus infeasible. Standard classical approaches for solving (1) in the general case suffer from the same problem, or they avoid it by not allowing for the optimization to include highly entangled states (see [19] for a review of some of these methods).
VQE instead parametrizes the set of state vectors by a parametrized quantum circuit $U(\theta)$ acting on an easily preparable initial state $|\psi_{0}\rangle$, leading to $|\psi(\theta)\rangle\coloneqq U(\theta)|\psi_{0}\rangle$, and measures the expectation of each observable $H_{j}$ with respect to $|\psi(\theta)\rangle$ in order to obtain an estimate $\widehat{E_{j}}$ of $\langle\psi(\theta)|H_{j}|\psi(\theta)\rangle$. Then the sum $\sum_{j}\widehat{E_{j}}$ is an estimate of the desired expectation $\langle\psi(\theta)|H|\psi(\theta)\rangle$. The estimates of $\langle\psi(\theta)|H|\psi(\theta)\rangle$ and its gradient can be fed to a classical optimization procedure, and this process iterated until convergence or a maximum number of iterations is reached. As a consequence of (3), this procedure in the ideal case produces an estimate of an upper bound on the ground-state energy $\lambda_{\min}(H)$.
@@ I.2 Main contribution: Dual-VQE method
The dual-VQE method involves 1) converting inequality constraints to equality constraints by means of positive semidefinite slack variables, 2) converting a constrained optimization into an unconstrained one by means of a penalty term in the objective function, 3) replacing positive semidefinite slack variables with scaled parametrized quantum states that are efficiently preparable on quantum computers, and 4) estimating various terms in the objective function by means of the destructive swap test [22] and sampling estimates of expectations of observables. Our companion paper [16] explores this general approach in a much broader context beyond the ground-state energy problem.
Although there have been various classical algorithms proposed for obtaining lower bounds on the ground-state energy [31, 32, 33, 4, 26], they all ultimately suffer from the exponential increase of the space on which $H$ acts, as the number of particles increases. Our approach is fundamentally different, involving a reformulation of the original optimization problem in terms of its SDP dual, and then bounding that quantity from below using a variational quantum algorithm. Our approach is also complementary to [10, Section VI-A], which does not take a variational approach, nor does it provably obtain a lower bound on the ground-state energy.
@@ I.3 Pretraining with matrix product states
A key issue that plagues variational quantum algorithms is the presence of barren plateaus in the training landscape [34, 21, 41], meaning that the gradient (or equivalently cost differences [2]) vanishes exponentially fast in the number of qubits. At their core, barren plateaus arise from a curse of dimensionality [13] stemming from the exponentially scaling size of the Hilbert space. On the other hand, standard techniques for provably avoiding barren plateaus do so by encoding the relevant dynamics in a smaller subspace making it effectively possible to classically simulate the parametrized quantum circuit [13].
@@ II Theoretical derivation
where $\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}$ is the SDP dual of (1). To see this, recall that the matrix inequality $A\geq B$ holds for Hermitian matrices $A$ and $B$ if ABA-B is a positive semidefinite matrix. Then consider that
@@
Now that we have recalled this standard result, let us focus on showing how to estimate the left-hand side of (4) on a quantum computer. Recalling that the matrix inequality $\eta I\leq H$ is equivalent to the existence of a positive semidefinite matrix $W$ such that $H-\eta I=W$, we conclude that
@@
Our next observation, related to observations used in [39], is that every positive semidefinite matrix $W\neq 0$ can be written as a scaled density matrix, i.e., $W=\nu\omega$, where $\nu=\operatorname{Tr}[W]$ and $\omega=W/\nu\in\mathcal{D}$, implying that
@@
where $c>0$ is a penalty parameter and $\left\|Z\right\|_{2}\coloneqq\sqrt{\operatorname{Tr}[Z^{{\dagger}}Z]}$ is the Hilbert–Schmidt norm of a matrix $Z$. Note that the equality in (13) holds whenever the penalty term is faithful (specifically, we invoke [8, Proposition 5.2.1] to justify this conclusion, setting $\lambda_{k}=0$ therein for all $k$). By faithfulness, we mean that the penalty term is equal to zero if and only if $H-\eta I=\nu\omega$. This is true for the Hilbert–Schmidt norm, and this particular norm has the convenient advantage that it can be estimated by quantum sampling procedures that we recall below.
The final step is to modify the optimization in (13) to be over a parametrized set $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of density matrices instead of being over all density matrices, which implies that
@@
The function $f^{\ast}(c)$ in (15) provides a lower bound on the ground-state energy of the Hamiltonian $H$ in the limit $c\to\infty$, i.e.,
@@
In the rest of the work, we use the term “unpenalized objective function” to refer to $\eta$, “penalty” to refer to the sum of all terms in (16) that are multiplied by the penalty parameter $c$, i.e.,
$\operatorname{Tr}[H^{2}]+\eta^{2}2^{n}+\nu^{2}\operatorname{Tr}[\omega(\theta)^{2}]\\ -2\eta\operatorname{Tr}[H]-2\nu\operatorname{Tr}[H\omega(\theta)]+2\eta\nu\,$, (18)
and “objective function” to refer to $f$ in (16) (i.e., the unpenalized objective function minus $c$ times the penalty).
@@ III Algorithm construction
The basic idea behind the dual-VQE algorithm is to follow the standard penalty method approximately [8, Section 5.2]: define a sequence $\left(f^{\ast}(c_{k})\right)_{k\in\mathbb{N}}$ of optimization problems where $\left(c_{k}\right)_{k\in\mathbb{N}}$ is a monotone increasing sequence of penalty parameters. We then solve the optimization for each $k$ and use the solution as the initial guess for the next iteration. Under the assumption that each optimization $f^{\ast}(c_{k})$ can be solved, the method is guaranteed to converge to the correct solution. In our simulations, we use a variety of update methods to change the value of $c$ as the optimization progresses. In some cases, we pick $c$ to be a high value and keep it constant throughout the optimization. In other cases, we begin with a smaller value, and ramp it up as the iteration count grows.
@@
As is common in Hamiltonian models relevant for physics and chemistry, the number of nonzero coefficients in the tuple $\left(\alpha_{\overrightarrow{x}}\right)_{\overrightarrow{x}}$ is polynomial inn~n, so that the objective function in (16) can be efficiently estimated. In this work, we focus on the transverse-field Ising model given in (32), which falls under this category.
@@
III.1 Parametrization
There are at least two general ways of parametrizing the set of mixed states, called the purification ansatz [17, 39, 20] and the convex combination ansatz [54, 29, 20, 46] (see also [16]). Focusing on the first one to start, it is based on the purification principle [12, 52, 53], in which every mixed state is realized as the marginal of a pure state on a larger system. For the purification ansatz, we simply start with a parametrized family $\left\{|\psi(\theta)\rangle_{RS}\right\}_{\theta\in\Theta}$ of pure states of a reference system $R$ and the system $S$ of interest, and then we obtain a parametrized family $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of mixed states via the partial trace $\omega(\theta)=\operatorname{Tr}_{R}[|\psi(\theta)\rangle\!\langle\psi(\theta)|_{RS}]$. We can estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ by means of the destructive swap test [22] (see, e.g., [3, Section 2.2] for a precise statement of the algorithm). We can estimate the gradient of the objective function in (17) by means of the parameter-shift rule [28, 36, 48], simply because
@@ III.1 Parametrization
$F$ is the unitary and Hermitian swap operator, and we have suppressed some system labels in (24) for brevity (see also [20, Eq. (C2)] in this context).
The convex combination ansatz is based on the fact that every mixed state can be written as a convex combination of pure orthogonal states. Following this idea, we obtain a parametrized family $\left\{\omega(\varphi,\gamma)\right\}_{\varphi,\gamma\in\Theta}$ by separately parametrizing the eigenvalues and eigenvectors as follows:
@@
where $\left(p_{\varphi}(x)\right)_{x}$ is a parametrized probability distribution, $U(\gamma)$ is a parametrized unitary, and $\left\{|x\rangle\right\}_{x}$ is the standard computational basis. The parametrized probability distribution $p_{\varphi}$ can be realized by a neural-network-based generative model [1, 6, 37] or a quantum circuit Born machine [5], while the parametrized unitary $U(\gamma)$ can be realized by a parametrized circuit.
Interestingly, although it is necessary to use a quantum computer to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ for the purification ansatz, it is not necessary to do so when using the convex combination ansatz in conjunction with a neural-network-based generative model: it suffices to employ a classical sampling approach, thus further reducing the quantum resource requirements for dual-VQE (see also [20, Eqs. (20)–(23)] in this context). This is because $\operatorname{Tr}[\left(\omega(\varphi,\gamma)\right)^{2}]=\sum_{x}p_{\varphi}^{2}(x)$, so that one can repeatedly sample from $p_{\varphi}$ and perform a collision test to estimate $\sum_{x}p_{\varphi}^{2}(x)$. Thus, for the convex combination ansatz used in the aforementioned way, one only needs a quantum computer to estimate the expectation
@@ III.2 Pretraining
Pretraining broadly refers to a set of computational techniques that are relatively inexpensive and kick start (or warm start) an optimization process, as opposed to a random start. In several domains, pretraining has been shown to be advantageous in reducing the training time and accelerating convergence [15, 35].
In this section, we describe the details of the particular pretraining process used in this work. The overall process consists of two phases: pretraining, followed by a translation into a quantum circuit. The pretraining is conducted using matrix product states (MPS) [18], and the final translation consists of a combination of two techniques, analytic decomposition (AD) and optimizing decomposition (OD), as proposed in [42, 43]. Important preliminaries on tensor networks and matrix product states can be found in Appendix A, and key ideas of the pretraining process, including an in-depth review of AD and OD, can be found in Appendix B.
Before delving into the specifics of the algorithm, we aim to establish the rationale behind pretraining using MPS. Our motivation for this approach stems from Refs. [44, 45], which explored solving the Variational Quantum Eigensolver (VQE) problem by pretraining it with MPS. The study demonstrated that parametrized quantum circuits for VQE exhibited significant performance improvements when used in conjunction with a pretrained initialization, in contrast to a random initialization. When applied to the dual-VQE problem, we find that pretraining provides a significant kick start, with the advantage increasing with maximum bond dimension of the MPS.